Please show work.

Analysis and Interpretation of Profitability Balance sheets and income statements for 3M Company follow. Consolidated Statements of Income Years ended December 31 (S millions) 2012 2011 2010 Net sales Operating expenses Cost of sales Selling, general and administrative expenses Research, development and related expenses $29,904 29,611 $226,662 15,685 15,693 13,831 6,102 6,1705,479 ,634 1,570 1434 23,421 23,433 20,744 Total operating expenses Operating income Interest expenses and income Interest expense Interest income 6,483 6,1785,918 186 (39) 147 201 (38) 163 6,351 6,031 5,755 1,592 4,511 $4,357 $4,163 78 $4,444 4,283 4,085 (39) 132 Total interest expense-net Income before income taxes Provision for income taxes Net income including noncontrolling interest Less: Net income attributable to noncontrolling interest Net income attributable to 3M 1,840 1,674 67 74
Answer #1

a) Nopat for 2012 is computed as [ 6483 - (1840 - ((171-39)*37%) = 4692

b)

2012 2011
Total Assets 33876 31616
Less Cash and cash equilvalent -2883 -2219
Less Investments -163 -155
Operting assets (A) 30830 29242
Total current liability 6200 5441
Less short term borrowing -1085 -682
Add Pension post retirement 3086 3972
Add other liabilties 1634 1857
Operating liability (B) 9835 10588
NOA (A-B) 20995 18654

c)

Formula 2012 Working
RNOA (NOPAT / NOA) 22.35% =4692 / 20995
NOPM (NOPAT / Sales) 15.69% =4692 / 29904
NOAT (Sales / NOA) 1.42 =29904 / 20995

d) Computation of NNO

2012 2011
Long term loan 4916 4484
Short term borrowing 1085 682
Less Cash and cash equivalent -2883 -2219
NNO 3118 2947

e) ROE = Net income attributable to 3m divided by Average of Total 3m company equity and Total equity

4444 / (17575 + 18040)/2 = 24.96%

f) Non operating return component = ROE - RNOA

= 24.96% - 22.35% = 2.61%

g) The best reference that can be drawn is ROE >RNOA imples that 3M is able to borrow money to fund operating assets that yield a return greater than its cost of debt
Answer #2

a) Calculation of NOPAT: NOPAT = EBIT x (1-Tax rate) - $6483 x (1-0.37) -$4084.29 b) Calculationof Net operating assets is asc) Calculate the return on net operating assets is as follows: Return on net operating assets = NOPAT/ Average net operating

d) Calcualtion of Net Operating obligations is as follows Particulars Cash and Cash Equivalents Invstments Current Marketable

As per HomeworkLib guidelines i have answered 4 sub parts as a part of time constraint.